Find the mean of the distribution shown below. x P(X) | 2 3 4 0.32 0.48 0.20 2.88 0 1.00 2.44 2.76
Answer #1

LO Mean = { x. Plu) = 2: PCA) + 3 P(3) +4. PM) = (23(0.32) + (3) 0:48) +410.20) 3 (0.64) +(1.44)+(0.go) = 2.88 Mean = 2.88 ]